OpenStudy (anonymous):

I really need help! solve the equation for y' = 0. y = 5x^2 + x -1 I dont know how to do this... I know it has to deal with the CHAIN RULE

OpenStudy (anonymous):

it should be -1/10. if you take the derivative of the function Y then y prime will be 10x+1 and if you set y prime equal to zero then x should be -1/10

OpenStudy (anonymous):

wait so after you get the derivative of y which is 10x+1 you do substitute zero???

OpenStudy (campbell_st):

actually it doesn't require the chain rule. all you need to do is 1. find the derivative 2. set the derivative equal to zero 3. solve the derivative for x

OpenStudy (campbell_st):

your derivative is correct y' = 10x + 1 now set it to zero 0 = 10x + 1 and solve for x

OpenStudy (anonymous):

y=5x^2+x-1 y'=dy/dx=10x+1=0 10x=-1,x=-1/10, y=5(-1/10)^2+(-1/10)-1=5/100-1/10 -1=(5-10-100)/100=-105/100=-21/20 x=-1/10, y=-21/20

OpenStudy (anonymous):

=∫▒〖logx xdx〗 ∫▒〖xlogx dx〗 =logxx^2/2-∫▒〖1/x x^2/2〗 dx =logx x^2/2-1/2 ∫▒xdx =logx x^2/2 -1/2 x^2/2 +c =logx *x^2/2 -x^2/4+c =∫▒〖logx xdx〗 ∫▒〖xlogx dx〗 =logxx^2/2-∫▒〖1/x x^2/2〗 dx =logx x^2/2-1/2 ∫▒xdx =logx x^2/2 -1/2 x^2/2 +c =logx *x^2/2 -x^2/4+c ∫▒〖logx xdx〗=∫▒〖xlogx dx〗 =logxx^2/2-∫▒〖1/x x^2/2〗 dx =logx x^2/2-1/2 ∫▒xdx =logx x^2/2 -1/2 x^2/2 +c =logx *x^2/2 -x^2/4+c